OpenStudy (anonymous):

simplify (x^2y^-1)^-3

OpenStudy (anonymous):

thats not one of my options on my answer sheet ?

OpenStudy (anonymous):

\[\huge (x^a)^b=x^{ab}\]

OpenStudy (anonymous):

you need to use the formula I gave you to solve the question

OpenStudy (anonymous):

Use the power rule for exponents. ChmE just gave it to you above.

OpenStudy (anonymous):

oh duh -_-

OpenStudy (anonymous):

k thanks let me try it

OpenStudy (anonymous):

yup thats what i got !

OpenStudy (anonymous):

@ktnguyen1 is incorrect

OpenStudy (anonymous):

idk why on my answer sheet its the oppisite way, does it matter?

OpenStudy (anonymous):

yes because they are wrong.

OpenStudy (anonymous):

\[ (x^2y^{-1})^{-3}=x^{-6}y^3=\frac{y^3}{x^6}\]

OpenStudy (anonymous):

so its y^3/x^6

OpenStudy (anonymous):

@AnimalAin is correct. look at how they applied the power rule for exponents

OpenStudy (anonymous):

OK, so you've got an answer. Do you understand how to do it?

OpenStudy (anonymous):

yup thanks guys!